OpenStudy (anonymous):

how do i do this? find f/g(x) for: f(x)= x / x+1 g(x)= x^2-1

OpenStudy (anonymous):

"f of g of x" so plug the g(x) into x in the f(x)

OpenStudy (anonymous):

so (x^2-1)/(x^2-1)+1

OpenStudy (anonymous):

find f/g(x) for each f(x) and g(x)

OpenStudy (anonymous):

is that f(x) DIVIDED by g(x)?

OpenStudy (anonymous):

f(x)= (x) / (x^2)+1 g(x)= (x^2)-1

OpenStudy (anonymous):

yes sorry (f/g)(x)

OpenStudy (anonymous):

my first answer was : (draw picture)

OpenStudy (anonymous):